MOA-II Obsevation Log

Return to index
Previous day Next day

DATE: 2009/03/17
OBSERVER: Paul
CONDITION: clear
COMMENT: Power failure. Pump on. Telescope position lost. Manual drive to zero search. Observing with small rotator offset problem.
Compressor gas pressure: 2.2 MPa
Expected Observation Time: START 20:04 END 05:50


Go to the Seeing Log

TARGETRUNTOTAL RUNTOTAL FIELDTOTAL EXPOSURE (sec)
GB17396-1740712744440
LMC-000
SMC-000
ALERT-000
GRB-000
TOTAL---12744440 (01h14m00s)


GB FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
2277811381211
Field13Field14Field15Field16Field17Field18Field19Field20Field21Field22Field23
37112221110


LMC FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
000000000000
Field13Field14Field15Field16
0000


SMC FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
000000000000


SPECIAL TARGET NAMEBR>
CategoryTargetNrun


CHECK SHEET

TimeStatusWeather
20:04waitingpumping
21:00waitingpumping
22:00waitingpumping
23:00waitingpumping . Zero search manually. Moatel recovery
00:00waitingZero search manually
01:00observingTest images. offset rotator instructions poor? -
02:00observingpartly cloudy
03:00observingpartly cloudy
04:00observingcloudy
05:00waitingcloudy
05:50waitingcloudy


CCD TEMPERATURE Previous day Next day




VACUUM Previous day Next day




TELESCOPE TEMPERATURE Previous day Next day




Return to index

Previous day Next day